# MoE-DiffuSeq: Enhancing Long-Document Diffusion Models with Sparse Attention and Mixture of Experts

### v1

@@ Abstract
In this paper, we present MoE-DiffuSeq, an innovative approach leveraging a Mixture of Experts (MoE) to enhance diffusion models for generating long documents. Traditional diffusion models, such as DiffuSeq, face significant computational inefficiencies and memory constraints when processing extended sequences. To overcome these challenges, we integrate a sparse attention mechanism with the MoE framework, facilitating efficient and scalable sequence generation. Our methodology incorporates a bespoke sparse attention framework designed to minimize computational complexity while preserving text quality and coherence. Additionally, we introduce a tailored soft absorbing state within the diffusion process to optimize the attention architecture, which accelerates sequence reconstruction and enhances precision. Empirical evaluations reveal that MoE-DiffuSeq substantially outperforms existing models in terms of training speed and sampling rates. These improvements are particularly beneficial for managing extended sequences, rendering our model apt for generating scientific articles, managing code repositories, and facilitating long dialogues. Benchmark comparisons further affirm that MoE-DiffuSeq enhances efficiency, speed, accuracy, and expressiveness, thereby advancing the capabilities of diffusion models for high-quality long-form text generation.
@@ Conclusion
Appendix A Preliminaries
Diffusion Models for Text Generation
Diffusion models have emerged as a robust framework for generative processes, particularly in the domain of text generation, where models like DiffuSeq exemplify their potential. These models operate through a two-phase approach: the forward diffusion process and the reverse diffusion process.
Forward Diffusion Process
During the forward diffusion phase, an initial data point, $x_{0}$, undergoes a transformation into a noise-dominated state through a sequence of intermediate states $x_{t}$ for $t=1,2,\ldots,T$, where $T$ signifies the total number of timesteps. The mathematical representation of this process is given by:
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$
where $\beta_{t}$ denotes the variance schedule, typically a small, incrementally increasing positive value. As $t$ increases, $x_{t}$ progressively approaches a purely noisy state, ultimately reaching $x_{T}\sim\mathcal{N}(0,I)$, symbolizing a standard Gaussian distribution.
Reverse Diffusion Process
The objective of the reverse diffusion process is to reconstruct the original data from its noisy counterpart. This reconstruction is facilitated by a learned denoising function $f_{\theta}(x_{t},t)$, which predicts $x_{t-1}$ from $x_{t}$:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};f_{\theta}(x_{t},t),\sigma_{t}^{2}I),$
where $\sigma_{t}$ represents a learned variance parameter. The process iteratively refines $x_{T}$, moving towards $x_{0}$ to effectively reverse the noise addition.
Handling Discrete States
In scenarios where the states are discrete, such as text tokens, each state $x_{t}$ is treated as a distinct entity. The forward diffusion might employ a categorical distribution, while the reverse process may involve transitions between discrete states, often facilitated by methods like masking or token substitution.
This dual-phase diffusion approach not only allows for the effective generation of high-quality text that captures intricate dependencies but also adapts well to the inherent complexities of structured and unstructured data. By leveraging the diffusion model framework within the DiffuSeq architecture, we can address significant challenges in text generation, particularly those associated with the management of long sequences and the preservation of semantic integrity.
Mixture of Experts (MoE) in Text Generation
The Mixture of Experts (MoE) paradigm revolutionizes large-scale neural network architectures by dynamically allocating computational resources among specialized sub-models, or ”experts,” to process each input. This modular approach significantly enhances scalability and adaptability, particularly when managing diverse and complex tasks.
Mathematical Representation
An MoE model is mathematically represented as:
$y=\sum_{i=1}^{N}g_{i}(x)E_{i}(x),$
where $N$ denotes the total number of experts, $g_{i}(x)$ is the gating function that allocates weights to the $i$-th expert $E_{i}$, based on the input $x$. The gating function is designed to produce a sparse output, ensuring that only a relevant subset of experts is activated for any given input, thereby optimizing computational resources.
Gating Mechanism
The gating mechanism, pivotal for the efficiency of MoE, is expressed as:
$g_{i}(x)=\frac{\exp(h_{i}(x))}{\sum_{j=1}^{N}\exp(h_{j}(x))},$
where $h_{i}(x)$ is a learnable scoring function that evaluates the suitability of each expert for the current input. This mechanism promotes sparsity and specialization by activating only the most relevant experts, thus reducing unnecessary computations.
Application in Long Text Generation
In the context of long text generation, MoE models are particularly advantageous. They dynamically allocate processing power to handle various segments of a text, adapting to the unique demands of each segment. This ability is crucial for generating coherent and contextually accurate extended texts, such as detailed scientific articles, elaborate narratives, and comprehensive reports.
By integrating MoE into the DiffuSeq framework, our approach—MoE-DiffuSeq—not only leverages the inherent strengths of diffusion models but also introduces a scalable method to manage the increased computational demands of processing lengthy documents. This integration ensures that our model can efficiently handle sequences of considerable length without sacrificing the quality of the generated content.
This dynamic allocation of computational resources, tailored to the specific requirements of different text segments, ensures that our generative process remains both efficient and effective, making MoE-DiffuSeq a cutting-edge solution in the realm of advanced text generation systems.
Sparse Attention Mechanism
Efficiently managing long text sequences presents a significant challenge in natural language processing. The sparse attention mechanism, exemplified by architectures like Longformer, directly addresses this by reducing the computational overhead traditionally associated with the self-attention mechanism of Transformer models. This reduction is achieved by altering the complexity from $O(n^{2})$ to $O(n\times w)$, where $n$ represents the sequence length and $w$ denotes the window size.
Mathematical Formulation
The sparse attention mechanism is mathematically defined as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$
where $Q$, $K$, and $V$ represent the query, key, and value matrices respectively, and $d_{k}$ is the dimensionality of the key vectors. In this configuration, each token computes attention weights only for a subset of tokens within a predefined window, significantly reducing the number of operations involved.
Implementation of Sparse Attention
Sparse attention is implemented as:
$\text{Sparse Attention}(Q,K,V)=\sum_{i=1}^{n}\sum_{j\in\mathcal{W}(i)}\text{softmax}\left(\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{k}}}\right)V_{j},$
where $\mathcal{W}(i)$ specifies the set of indices within the attention window of the token $i$. This method efficiently limits the scope of attention, focusing computational resources where they are most needed.
Enhancements for Long Context
To capture dependencies over longer distances without increasing computational demands, dilated attention is utilized:
$\mathcal{W}(i)=\{i-k\cdot d,\ldots,i+k\cdot d\},$
where $d$ represents the dilation factor, enhancing the model’s ability to integrate information over larger textual spans without a corresponding increase in the window size.
Additionally, global attention tokens are incorporated to provide context integration across the entire sequence. This feature is particularly valuable in tasks that require an overarching understanding of the full text, such as in document summarization or complex question-answering scenarios.
Advantages of Sparse Attention
Integrating sparse attention mechanisms allows models to efficiently process extensive documents while ensuring the preservation of text quality and coherence. This approach not only mitigates the quadratic scaling issue but also enhances the model’s capability to generate contextually rich and coherent long-form text. By focusing computational resources judiciously, sparse attention mechanisms facilitate deeper and more effective text analysis and generation, especially in applications involving substantial textual data.

### v2

@@ Abstract
We propose MoE-DiffuSeq, a diffusion-based framework for efficient long-form text generation that integrates sparse attention with a Mixture-of-Experts (MoE) architecture. Existing sequence diffusion models suffer from prohibitive computational and memory costs when scaling to long documents, largely due to dense attention and slow iterative reconstruction. MoE-DiffuSeq addresses these limitations by combining expert routing with a tailored sparse attention mechanism, substantially reducing attention complexity while preserving global coherence and textual fidelity. In addition, we introduce a soft absorbing state within the diffusion process that reshapes attention dynamics during denoising, enabling faster sequence reconstruction and more precise token refinement. This design accelerates both training and sampling without sacrificing generation quality. Extensive experiments on long-document benchmarks demonstrate that MoE-DiffuSeq consistently outperforms prior diffusion-based and sparse-attention baselines in training efficiency, inference speed, and generation quality. Our approach is particularly effective for long-context applications such as scientific document generation, large-scale code synthesis, and extended dialogue modeling, establishing a scalable and expressive solution for diffusion-based long-form text generation.